speed of a dropped object. If we neglect the resistnace of air, the (downward) speed of an object t seconds after being dropped is given by v=32t feet per second. How many seconds does it take an object to reach a downward speed of 100 miles per hour? recall that there are 5280 feet in one mile, since you will need to express 100 mph in feet per second.
Answer by fractalier(2101)

Okay...we can convert mph to fps via the relationship
60 miles per hour = 88 feet per second so that
100 miles per hour = 146 2/3 feet per second.
Thus we solve for t and get
t = v/32
t = (146 2/3) / 32
t = 55/12 sec = 4 7/12 sec